# Intrinsic Diophantine approximation: a solution to Mahler’s problem

### v1

@@ Abstract.
In the 80’s Mahler suggested to study the Diophantine approximation by rationals lying in the middle-third Cantor set $K_{1/3}$. In this article we address this question by proving that, for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+},$
$\dim_{H}\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }\frac{p}{q}\in\mathbb{Q}\cap K_{1/3}\right\}=\frac{\dim_{H}K_{1/3}}{\max\left\{1,\delta_{\psi}\right\}},$
where $\delta_{\psi}$ is the shrinking rate of $\psi.$ This result will be derived out a more general theorem regarding homogeneous self-similar rational IFS’s. In addition, we establish several results regarding the intrinsic Diophantine approximation associated with inhomogeneous rational IFS’s.
@@ 1. Introduction
Among these results, Benard, He and Zhang [3] showed that if $\mu$ is a self-similar measure on $\mathbb{R}$ (Definition 2.3), then for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$ one has (where “f.i.m.” stands for “for infinitely many”)
@@
$\dim_{H}\Big\{x\in K_{1/3}:\ |x-\tfrac{p}{q}|\leq q^{-\delta}\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{1/3}\Big\}$
@@
Here, the first term corresponds to the intrinsic contribution (obtained by considering the approximation by rationals lying in the middle-third Cantor set), while the second corresponds to the extrinsic one. In this direction, Chow, Yu and Varjú [9] established that, in the case of the middle-seventh Cantor set, there exists $\delta_{0}>2$ such that the above formula holds for any $2\leq\delta\leq\delta_{0}$. This result was later generalized by Chen (using techniques from [3]) to any self-similar set satisfying the open set condition [8].
Regarding intrinsic approximation, namely approximation by rationals lying inside a given self-similar set, it is worth mentioning that it is not generic for a self-similar set to contain rational numbers. Indeed, given a homogeneous self-similar set $K\subset\mathbb{R}$, it is easily checked that for every $y\in\mathbb{R}$, $y+K$ is also the attractor of a homogeneous self-similar IFS. Moreover, one can show that
@@
A natural class of self-similar sets known to contain rational numbers consists of those defined by rational parameters. The first results in this setting were obtained by Beresnevich and Velani [4], and refined by Levesley, Salp and Velani [20], who provided a complete description of the metric theory of $\psi$-approximable points in base 3 in $K_{1/3}$. The problem was also studied, among others, by Fishman and Simmons [18], Tan, Wang and Wu [24], and Baker [1]. Writing $h(r)$ the height of $r\in\mathbb{Q},$ that is, $h(r)=q$ if $r=\frac{p}{q}$ with $p\wedge q=1$, in [18] and [24], the authors proposed the following conjecture (expressed differently), extending that of Bugeaud and Durand: for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$, writing $\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q},$ [^1] A careful reader will notice that the set investigated by Bugeaud and Durand does not involve the height $h(r)$. It as small exercise to show that, for any non increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$ $\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }(p,q)\in\mathbb{Z}\times\mathbb{N}\text{ s.t. }\frac{p}{q}\in K_{T}\right\}=\left\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\text{ f.i.m. }r\in\mathbb{Q}\cap K_{T}\right\}\bigcup(\mathbb{Q}\cap K_{T})$. Thus, the choice of the definition of the approximating sets does not change the dimension results.
$\dim_{H}\Big\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{1/3}\Big\}=\frac{\dim_{H}K_{1/3}}{\max\Big\{1,\ \delta_{\psi}\Big\}}.$
We confirm this conjecture by proving the following result.
@@
Let $b\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$ be integers, and set
@@
Let $K_{T}$ be the attractor of $T$ (see Definition 2.3). If $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, writing
$\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q},$
we have
$\dim_{H}E_{\psi,T}:=\dim_{H}\Big\{x\in K_{T}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}}.$
Notice that, in the above formula, $\frac{\dim_{H}K_{T}}{\max\left\{1,\delta_{\psi}\right\}}=\min\left\{\dim_{H}K_{T},\frac{\dim_{H}K_{T}}{\delta_{\psi}}\right\}.$ Moreover, it is a routine check to verify that
@@
In contrast, by Khintchine’s theorem, when $K_{T}$ has non empty interior, one has, for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$
$\dim_{H}E_{\psi,T}=\min\left\{1,s_{\psi}\right\},$
@@
$s_{\psi}=\inf\left\{s\geq 0:\ \sum_{n\geq 1}n\psi(n)^{s}<+\infty\right\}.$
@@
Let $K_{T}$ be the attractor of $T$, $\delta\geq 1$ and write $\psi_{\delta}:q\mapsto q^{-\delta}.$ Then
@@
The second purpose of this article is to provide more general results for inhomogeneous rational IFSs. Given integers $m\geq 2$, $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, for $1\leq i\leq m$ define, for any $x\in\mathbb{R},$
@@
let $T=\{f_{1},\dots,f_{m}\}$ and denote by $K_{T}$ its attractor. We recall the following fact.
Lemma 1.3 ([18, 1]).
@@
$r=\pi\Big((i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty})\Big),$
where $\pi$ is the canonical projection associated with $T$.
@@
$\dim_{H}E_{\psi,T,\text{int}}=\frac{\dim_{H}K_{T}}{\max\Big\{1,\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}\Big\}}.$
It is worth mentioning that we do not require any separation condition on $T$. In particular, it is not known in general whether $\mathcal{H}^{\dim_{H}K_{T}}(K_{T})>0$ for IFSs of the form (1). Moreover, this result follows from a more general theorem concerning approximation by ultimately periodic sequences on self-similar fractals (see Theorem 4.1).
@@
$\dim_{H}\Big\{x\in K_{T}:\ |x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{T}\Big\}\geq\frac{\dim_{H}K_{T}}{\max\left\{1,\liminf_{n\to+\infty}\frac{-\log\psi(n)}{\log n}\right\}}.$
2. Notations, recalls and main statements
@@
Given $E\subset\mathbb{R}^{d}$, $\dim_{H}(E)$ and $\dim_{P}(E)$ denote respectively the Hausdorff and the packing dimension of $E$.
@@ 2.1. Recalls on geometric measure theory and fractal geometry
For $t\in(0,+\infty]$, $s\geq 0$ and $\zeta(x)=x^{s}$, we write $\mathcal{H}^{\zeta}_{t}(E)=\mathcal{H}^{s}_{t}(E)$ and $\mathcal{H}^{\zeta}(E)=\mathcal{H}^{s}(E)$, which are respectively called the $s$-dimensional Hausdorff outer measure at scale $t$ and the $s$-dimensional Hausdorff measure. Thus,
(3) $\mathcal{H}^{s}_{t}(E)=\inf\left\{\sum_{n\in\mathbb{N}}|B_{n}|^{s}:\ |B_{n}|\leq t,\ B_{n}\text{ closed balls and }E\subset\bigcup_{n\in\mathbb{N}}B_{n}\right\}.$
The quantity $\mathcal{H}^{s}_{\infty}(E)$ is called the $s$-dimensional Hausdorff content of $E$.
@@
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$. For $x\in\mathrm{supp}(\mu)$, the lower and upper local dimensions of $\mu$ at $x$ are defined by
$\underline{\dim}_{\rm loc}(\mu,x)=\liminf_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r},\qquad\overline{\dim}_{\rm loc}(\mu,x)=\limsup_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r}.$
The lower Hausdorff dimension and upper packing dimension of $\mu$ are respectively defined by
(4) $\dim_{H}(\mu)=\mathrm{ess\,inf}_{\mu}(\underline{\dim}_{\rm loc}(\mu,x)),\qquad\overline{\dim}_{P}(\mu)=\mathrm{ess\,sup}_{\mu}(\overline{\dim}_{\rm loc}(\mu,x)).$
@@
$\dim_{H}(\mu)=\inf\{\dim_{H}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\ \mu(E)>0\},$
$\overline{\dim}_{P}(\mu)=\inf\{\dim_{P}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\ \mu(E)=1\}.$
When $\dim_{H}(\mu)=\overline{\dim}_{P}(\mu)$, this common value is denoted by $\dim_{H}(\mu)$ and $\mu$ is said to be exact dimensional.
2.2. Self-similar iterated function systems

### v3

# Intrinsic Diophantine approximation by rationals of height with a bounded number of distinct prime factors

@@ Abstract.
In the 80’s Mahler suggested to study the Diophantine approximation by rationals lying in the middle-third Cantor set $K_{1/3}$. In this article, writing, for every $N\in\mathbb{N},$ $\mathcal{P}_{N}$ the set of numbers with less than $N$ distinct prime factors, we show that, for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+},$
$\dim_{H}\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }\frac{p}{q}\in\mathbb{Q}\cap K_{1/3},\ q\in\mathcal{P}_{N}\right\}=\frac{\dim_{H}K_{1/3}}{\max\left\{1,\delta_{\psi}\right\}},$
where $\dim_{H}\cdot$ denotes the Hausdorff dimension and $\delta_{\psi}$ is the shrinking rate of $\psi.$ This result will be derived out a more general theorem regarding homogeneous self-similar rational IFS’s. In addition, we establish several results regarding the intrinsic Diophantine approximation associated with inhomogeneous rational IFS’s.
@@ 1. Introduction
Among these results, Bénard, He and Zhang [3] showed that if $\mu$ is a self-similar measure on $\mathbb{R}$ (Definition 2.3), then for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$ one has (where “f.i.m.” stands for “for infinitely many”)
@@
$\dim_{H}\Big\{x\in K_{1/3}:\ |x-\tfrac{p}{q}|\leq q^{-\delta}\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\Big\}$
@@
Here, the first term corresponds to the intrinsic contribution (obtained by considering the approximation by rationals lying in the middle-third Cantor set), while the second corresponds to the extrinsic one. In this direction, Chow,Varjú and Yu [10] established that, in the case of the middle-seventh Cantor set, there exists $\delta_{0}>2$ such that the above formula holds for any $2\leq\delta\leq\delta_{0}$. This result was later generalized by Chen [9] and He and Liao [20] to self-similar sets satisfying the open set condition.
Regarding intrinsic approximation, namely approximation by rationals lying inside a given self-similar set, it is worth mentioning that it is not generic for a self-similar set to contain rational numbers. Indeed, given an homogeneous self-similar set $K\subset\mathbb{R}$, it is easily checked that for every $y\in\mathbb{R}$, $y+K$ is also the attractor of an homogeneous self-similar IFS. Moreover, one can show that
@@
A natural class of self-similar sets known to contain rational numbers consists of those defined by rational parameters. The first results in this setting were obtained by Beresnevich and Velani [4], and refined by Levesley, Salp and Velani [23], who provided a complete description of the metric theory of $\psi$-approximable points in base 3 in $K_{1/3}$. The problem was also studied, among others, by Fishman and Simmons [19], Tan, Wang and Wu [27], and Baker [1]. Writing $h(r)$ the height of $r\in\mathbb{Q},$ that is, $h(r)=q$ if $r=\frac{p}{q}$ with $p\wedge q=1$, in [19] and [27], the authors proposed the following conjecture (expressed differently), extending that of Bugeaud and Durand [^1] A careful reader will notice that the set investigated by Bugeaud and Durand does not involve the height $h(r)$. It as an exercise to show that, for any non increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$ $\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }(p,q)\in\mathbb{Z}\times\mathbb{N}\text{ s.t. }\frac{p}{q}\in K_{T}\right\}=\left\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\text{ f.i.m. }r\in\mathbb{Q}\cap K_{T}\right\}\bigcup(\mathbb{Q}\cap K_{T})$. Thus, the choice of the definition of the approximating sets does not change the dimension results.: for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$, writing
(1) $\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q},$
(2) $\dim_{H}\Big\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{1/3}\Big\}=\frac{\dim_{H}K_{1/3}}{\max\Big\{1,\ \delta_{\psi}\Big\}}.$
Given an integer $n\in\mathbb{N}$ who’s decomposition in prime numbers is $n=\prod_{i=1}^{k_{n}}p_{i}^{\alpha_{i}},$ we write
$\omega(n)=k_{n}$
and given $N\in\mathbb{N},$ we set
$\mathcal{P}_{N}=\left\{q\in\mathbb{N}:\ \omega(q)\leq N\right\}.$
In this article we make a step toward this conjecture by proving the following.
@@
Let $b\geq 2$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$ be integers and define
@@
Let $K_{T}$ be the attractor of $T$ (see Definition 2.3). For any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, write
$E_{\psi,T,N}=\Big\{x\in K_{T}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T},\ h(r)\in\mathcal{P}_{N}\Big\}.$
If $K$ has empty interior, then
  • (1)
    for every $N\geq\omega(b)+\omega(b-1),$
    $\dim_{H}E_{\psi,T,N}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}},$
    where $\delta_{\psi}$ is defined as in (1).
  • (2)
    if there exists $p\in\mathbb{Z},k\in\mathbb{N}$ such that $\frac{p}{b^{k}}\in K_{T},$ then item $(1)$ holds for every $N\geq\omega(b).$
Note the last item holds as soon as there exists $1\leq i\leq m$ such that $p_{i}=0\text{ or }b-1.$ In addition, when $b$ is a power of a prime, $\omega(b)=1.$ Hence, for the middle third Cantor set, item $(1)$ holds for every $N\geq 1.$
Remark 1.2.
For a self-similar IFS $T$ of the form $T=\Big\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\Big\}_{1\leq i\leq m}$, writing $K_{T}$ its attractor, the three following are equivalent:
  • (1)
    $\mathcal{L}(K_{T})>0,$
  • (2)
    $K_{T}$ has non empty interior,
  • (3)
    $\dim_{H}K_{T}=1.$
We refer to the last subsection of the appendix for a proof of this remark.
In the above formula, $\frac{\dim_{H}K_{T}}{\max\left\{1,\delta_{\psi}\right\}}=\min\left\{\dim_{H}K_{T},\frac{\dim_{H}K_{T}}{\delta_{\psi}}\right\}.$ Moreover, it is a routine check to verify that
@@
Denote $\phi$ is the Euler mapping, that is
$\phi(\cdot)=\#\left\{t\leq n,\ t\wedge n=1\right\}$
and notice that for any $n\in\mathcal{P}_{N},$
$\phi(n)\geq n\prod_{i=2}^{N+1}(1-\frac{1}{i})=C_{N}n.$
In contrast with Theorem 1.1, when $K_{T}$ has non empty interior, for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$ a direct application of [22] and [4] yields
$\dim_{H}E_{\psi,T,N}=\min\left\{1,s_{\psi}\right\},$
@@
$s_{\psi}=\inf\left\{s\geq 0:\ \sum_{n\in\mathcal{P}_{N}}n\psi(n)^{s}<+\infty\right\}.$
@@
Let $K_{T}$ be the attractor of $T$, $\delta\geq 1$ and $\psi_{\delta}:q\mapsto q^{-\delta}.$ Then, for any $N\geq\omega(b)+\omega(b-1),$ [^2] In the case where $K_{T}$ has non empty interior, the formula for $\dim_{H}E_{\psi_{\delta},T,N}$ is easily obtained from the expression of $s_{\psi}$ together with the fact that, for every large enough $x\in\mathbb{N},$ $\left\{n\leq x:\ \omega(n)\leq N\right\}\geq\#\left\{n\leq x:n\text{ is prime}\right\}\geq\frac{x}{2\log x}.$
@@
Regarding the original problem, we make a connection with the validity of the dimension formula proposed in (2) and a natural conjecture in number theory. Given $q,b\in\mathbb{N}$ such that $q\wedge b=1,$ we write $O_{q}(b)$ the order of $b$ in $\mathbb{Z}_{q}^{*},$ that is,
$O_{q}(b)=\min\left\{k\geq 1:\ b^{k}=1\mod q\right\}.$
Conjecture 1.4.
Let $b\geq 2$ be an integer and let $f:\mathbb{N}\to\mathbb{R}_{+}$ be a non-increasing mapping such that $f(q)\to 0.$ Then
$\lim_{x\to+\infty}\frac{\log\#\left\{1\leq q\leq x:\ q\wedge b=1,\ O_{q}(b)\leq q^{f(q)}\right\}}{\log x}=0.$
Theorem 1.5.
Let $b\geq 2$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$ be integers and define
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\Big\}_{1\leq i\leq m}.$
Let $K_{T}$ be the attractor of $T$. For any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, write
$E_{\psi,T}=\Big\{x\in K_{T}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\}.$
Assume that Conjecture 1.4 holds true and that $K_{T}$ has empty interior, then
$\dim_{H}E_{\psi,T}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}},$
where $\delta_{\psi}$ is defined as in (1).
If true, Conjecture 1.4 is probably difficult to prove. Nonetheless, we motivate this conjecture by proving a weaker version of it (which we use in the proof of Theorem 1.1).
Theorem 1.6.
Let $b\geq 2$ be an integer and let $f:\mathbb{N}\to\mathbb{R}_{+}$ be a non increasing mapping with $f(q)\to 0$. Then, for any $N\in\mathbb{N},$
$\lim_{x\to+\infty}\frac{\log\#\left\{1\leq q\leq x:\ q\wedge b=1,\ \omega(q)\leq N\ O_{q}(b)\leq q^{f(q)}\right\}}{\log x}=0.$
The second purpose of this article is to provide more general results for inhomogeneous rational IFSs. Given integers $m\geq 2$, $q_{1},\dots,q_{m}\geq 2$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, for $1\leq i\leq m$ define, for any $x\in\mathbb{R},$
@@
let $T=\{f_{1},\dots,f_{m}\}$ and denote by $K_{T}$ its attractor. We recall the following result, proved, which is due to Fishman and Simmons and Baker.
Proposition 1.7 ([19, 1]).
@@
[…]
where $\pi$ is the canonical projection associated with $T$, define for any $(j_{n})_{n\in\mathbb{N}}\in\left\{1,...,m\right\}^{\mathbb{N}}$ by
$\pi\Big((j_{n})_{n\in\mathbb{N}}\Big)=\lim_{n\to+\infty}f_{j_{1}}\circ...\circ f_{j_{n}}(0).$
As Proposition 1.7 was only proved in the case where $T$ satisfies the open set condition in [19] and mentioned to hold in [1] , for the sake of completeness, Proposition 1.7 is proved in the first subsection of the Appendix (see Subsubsection 6.1.2).
@@
$\dim_{H}E_{\psi,T,\text{int}}=\frac{\dim_{H}K_{T}}{\max\Big\{1,\delta_{\psi}\Big\}}.$
It is worth mentioning that it is not known in general whether $\mathcal{H}^{\dim_{H}K_{T}}(K_{T})>0$ for IFSs of the form (3). Moreover, this result follows from a more general theorem concerning approximation by ultimately periodic sequences on self-similar fractals (see Theorem 4.1).
@@
$\dim_{H}\Big\{x\in K_{T}:\ |x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{T}\Big\}\geq\frac{\dim_{H}K_{T}}{\max\left\{1,\delta_{\psi}\right\}}.$
In the first section, we recall some background on geometric measure theory, self-similar IFS’s and the separation properties they satisfy when the parameters defining these IFS’s are rationals. Since Theorem 1.8 provides the appropriate lower bound for Theorem 1.1 (see Corollary 1.9), the proof of Theorem 1.1 is split into two parts. Section 3 establishes the “upper-bound” part of Theorem 1.1, while Section 4 proves Theorem 1.8. Finally, Section 5 presents some open problems and perspectives related to the topic.
2. Notations and recalls
@@
Given two integers $n,m$ one will write $n|m$ if there exists $k\in\mathbb{Z}$ such that $m=nk$.
Given $f:\mathbb{N}\to\mathbb{R}_{+,}$ we will write $f(q)=o_{q\to+\infty}(1)$ if $\lim_{q\to+\infty}\frac{f(q)}{q}=0.$
@@ 2.1. Recalls on geometric measure theory and fractal geometry
For $t\in(0,+\infty]$, $s\geq 0$ and $\zeta(x)=x^{s}$, we write $\mathcal{H}^{\zeta}_{t}(E)=\mathcal{H}^{s}_{t}(E)$ and $\mathcal{H}^{\zeta}(E)=\mathcal{H}^{s}(E)$, which are respectively called the $s$-dimensional Hausdorff outer measure at scale $t$ and the $s$-dimensional Hausdorff measure.
@@
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$. For $x\in\mathrm{supp}(\mu)$, the lower dimension of $\mu$ at $x$ is defined by
$\underline{\dim}_{\rm loc}(\mu,x)=\liminf_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r}.$
The Hausdorff dimension of $\mu$ is defined by
(5) $\dim_{H}(\mu)=\mathrm{ess\,inf}_{\mu}(\underline{\dim}_{\rm loc}(\mu,x)).$
@@
$\dim_{H}(\mu)=\inf\{\dim_{H}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\ \mu(E)>0\}.$
2.2. Some geometric properties of self-similar iterated function systems